OpenStudy (anonymous):

for graphs y=a^x and y= log a (x), where a>0, what is the value of a that will cause these graphs to touch but not pass through each other?

OpenStudy (anonymous):

try plotting

OpenStudy (anonymous):

u have y = a^x and x = a^y, right?